# SMI: An Information-Theoretic Metric for Predicting Model Knowledge Solely from Pre-Training Signals

### v3

@@ Abstract
The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent. Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets. In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task indicative of a model’s internal knowledge. We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance. In response to these challenges, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models. Subsequently, we develop a multi-template QA evaluation framework incorporating paraphrased question variants. Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring any additional training. The experimental results demonstrate that SMI outperforms co-occurrence-based baselines, achieving R² > 0.75 on models with over one billion parameters. Theoretical analysis further reveals the marginal benefits of scaling model size and optimizing data, indicating that the upper limit of specific QA task accuracy is approximately 80%. Our project is available at https://github.com/yuhui1038/SMI.
@@ 1 Introduction
The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" (OpenAI, 2023). However, the report provides little detail on the methodology behind this prediction approach.
This technique is critical for the pre-training of large language models (LLMs). On the one hand, the cost of a single pre-training run can reach hundreds of millions of dollars, making re-pre-training a
@@
highly challenging endeavor (Jiang et al., 2024; Yang et al., 2024; Dubey et al., 2024). The prediction technique, which allows for accurate performance forecasting on certain tasks based only on the pre-training signals, helps mitigate the resource waste. On the other hand, it can guide the construction of pre-training data that are more aligned with target tasks, optimizing resource usage (Gururangan et al., 2020; Longpre et al., 2024).
@@
In this paper, we aim to predict performance in closed-book QA, a vital downstream task indicative of a model’s internal knowledge (Roberts et al., 2020). This setting is particularly well aligned with the nature of pre-training data and best reflects a model’s ability to answer fact-based questions. Meanwhile, we encounter three major challenges (Figure 1):
@@
Challenge1: Limited access to and understanding of pre-training corpora. Currently, few open-source foundational LLMs fully disclose their pre-training datasets. Initiating pre-training from scratch is a costly endeavor, requiring the collection of extensive data and substantial computational resources. For instance, Meta’s technical report states that pre-training the 405B-parameter Llama 3 model used 15.6 trillion tokens and 16,000 H100 GPUs, with estimated costs exceeding 100 million dollars (Dubey et al., 2024).
@@
Challenge2: Limitations of current evaluation methods for pre-trained models. Given the characteristics of closed-book QA tasks, we can evaluate a model’s accuracy on these tasks to assess the factual knowledge it has internalized during pretraining (Roberts et al., 2020; Wang et al., 2021). However, most evaluation methods face challenges, such as being overly sensitive to specific in-context examples and having coarse granularity in test data segmentation (Min et al., 2022; Srivastava et al., 2024). These issues make it difficult to accurately assess whether the pre-trained model has effectively retained specific knowledge.
@@
Challenge3: Limitations of frequency-based metrics in predicting model performance. The model’s success on the target task relies on its ability to acquire world knowledge during pre-training, a process heavily shaped by the data (Allen-Zhu and Li, 2024a; Wang et al., 2024). Previous research suggests that models tend to acquire knowledge more effectively when it is frequently encountered within the pre-training data (Kandpal et al., 2023; Mallen et al., 2023; Allen-Zhu and Li, 2024b). However, these frequency-based methods does not take into account the specificity of knowledge, and thus only demonstrates general trends. For example, despite similar frequencies of occurrence, the model may struggle to pinpoint the location of Apple’s headquarters due to the pervasive nature of the subject. In contrast, identifying NVIDIA’s headquarters might be relatively straightforward. Furthermore, frequency-based methods do not consider the impact of model size.
@@
Second, inspired by the evaluation methodology for fine-tuned LLMs (Ye et al., 2024), we develop a robust multimodal evaluation approach. For each category of QA tasks, we design 20 semantically equivalent but syntactically diverse evaluation templates, enabling a more accurate and reliable assessment of the model’s knowledge retention.
Finally, we introduce SMI, an information-theoretic metric that linearly associates pre-training data characteristics, model size, and QA accuracy without requiring additional training. Experimental results on 24 pre-trained models indicate that the SMI metric effectively predicts a model’s knowledge retention. Across various model sizes, ranging from 14 million to 176 billion parameters, SMI significantly outperforms the co-occurrence-based baseline. For models with more than 1 billion parameters, the coefficient of determination (R²) between SMI and QA accuracy exceeds 0.75, indicating a strong linear relationship.
Moreover, the comparative analysis of experimental results from different models further reveals that both scaling model size and optimizing data present diminishing returns. As the model size increases, its impact on performance in QA tasks becomes progressively smaller. Meanwhile, even when the SMI metric reaches its maximum on the largest models within each families, the accuracy of the QA tasks cannot surpass a ceiling of approximately 80%.
Overall, our contributions are threefold:
@@
    Experimental results across diverse models reveal that both scaling model size and optimizing data yield diminishing returns. Even under optimal conditions, the upper bound of accuracy on the evaluated QA tasks remains around 80%.
@@ 2 Related Work
The characteristics of pre-training data are crucial for the knowledge retention and downstream task performance of LLMs. Numerous studies highlight the significance of repetition frequency in this process. For instance, Carlini et al. (2023) identified a logarithmic relationship between repetition frequency and memory effects, while Chowdhery et al. (2023) demonstrated that sequences repeated over 500 times are completed by models with an accuracy exceeding 40%. Similarly, Ju et al. (2024) investigated the influence of data frequency on multi-hop reasoning, uncovering the presence of “factual shortcuts”. Allen-Zhu and Li (2024b) proposed that exposing knowledge 1000 times can achieve a storage capacity of two bits per parameter in LLMs. Additionally, Razeghi et al. (2022) and Yadlowsky et al. (2023) demonstrated that lower-order co-occurrences in pre-training data enhance numerical reasoning, while McCoy et al. (2023) associated the prevalence of tasks in pre-training data with better performance on tasks like ROT13.
@@
Furthermore, the model size significantly influences the knowledge retention and capabilities of LLMs. Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern (Kaplan et al., 2020).  Kandpal et al. (2023) evaluate the BLOOM model families (Scao et al., 2022) on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters. Additionally, the memory capacity of a model closely correlates with its size (Li et al., 2024).
@@
Knowledge triples (subject, relation, object) serve as an effective tool for evaluating the knowledge retention of LLMs. Petroni et al. (2019) employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks. Subsequently, He et al. (2024); Ju et al. (2024) further explore the application of knowledge triples in reasoning tasks. Moreover, Wang et al. (2024) and Allen-Zhu and Li (2024a) use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
@@ 3 Methodology
To address the three challenges outlined in Section 1, we propose a method to predict model knowledge solely from pre-training signals. We begin with prior work and explore the use of the co-occurrence metric. However, the co-occurrence metric do not account for the specificity of knowledge. Inspired by information theory, we propose the Mutual Information (MI) metric that simultaneously considers both the frequency and specificity of knowledge, significantly enhancing prediction outcomes. Nevertheless, the MI metric does not consider the impact of model size, which has been shown to correlate closely with model knowledge retention. Finally, we propose the SMI metric, a composite representation of pre-training data characteristics and model size.
@@ 3.1 Starting from the Co-occurrence Metric
Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns (Kandpal et al., 2023; Mallen et al., 2023; Allen-Zhu and Li, 2024b). Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach (Kandpal et al., 2023). There is a linear relationship between QA accuracy and the logarithm of the number of documents (Kandpal et al., 2023). The logarithm of the document count serves as the co-occurrence metric, which constitutes our baseline.
@@
In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object (Wang et al., 2024). For a given LLM, denoted as $F$, we define $F$ as mastering the knowledge triple $t$, thus demonstrating competence in the corresponding QA task, as follows:
@@
Evaluating the knowledge retention of LLMs presents challenges due to the diverse range of queries formed by the subject and relation of knowledge triples. Inspired by the evaluation method for fine-tuned LLMs (Ye et al., 2024), we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
@@
On the other hand, using the multi-template evaluation method, we obtain the model’s accuracy on QA tasks. We fit a linear equation between the co-occurrence metric and QA accuracy, and then calculate R² and Mean Square Error (MSE) to assess the strength and stability of their linear relationship.
@@ 3.2 Considering the Knowledge Specificity to Obtain the MI Metric
The co-occurrence metric does not account for the specificity of knowledge. When entities of knowledge are common phrases that frequently appear in pre-training data, their specificity is low, making the knowledge more difficult to retain. Conversely, if the entities appear infrequently in pre-training data, their specificity is high, and the knowledge is easier to retain. For example, the address of NVIDIA’s headquarters is easier to remember than that of Apple’s headquarters, because the word "Apple" occurs more frequently across diverse contexts. MI in information theory considers both the co-occurrence of two variables and their individual specificities (Shannon, 1948). Therefore, we aim to use MI to improve the co-occurrence metric.
@@
Unlike the MI in information theory, which considers all possible values of random variables, the MI metric within pre-training data considers only those instances in which the relevant entities actually appear. Given pre-training data $\mathbf{P}$ consisting of $\mathbf{N}$ paragraphs, a knowledge triple $(s,r,o)$, and corresponding $P(s)$, $P(o)$, and $P(s,o)$, the MI metric between $s$ and $o$ is defined as:
@@
In the above formula, $P(s)$ and $P(o)$ penalize the number of documents containing the subject and object, respectively, to emphasize the specificity of $s$ and $o$. The term $I(s,o)$ represents the amount of shared information between $s$ and $o$, and it can also indicate the reduction in the amount of information about $o$ given $s$. During inference with LLMs, this amount of information refers to the probability of the model generating the object given the subject. Considering the uneven distribution of $I(s,o)$, we take its logarithm to obtain the MI metric (Figure 2).
@@ 3.3 Considering the Model Size to Obtain the Size-dependent MI metric
The MI metric does not account for the impact of model size. Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger (Kaplan et al., 2020). In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters (Kandpal et al., 2023). Therefore, we propose to enhance the MI metric by incorporating model size $\Phi$, measured in billions of parameters.
@@
4 Experiments
First, we clean and enhance an open-source Wikipedia knowledge triple dataset to make it suitable for inference scenarios using pre-trained models. Next, we conduct retrieval analysis on five large-scale open-source pre-trained corpora. Finally, we perform experimental analysis on 21 open-source pre-trained models and 3 models that we pre-train ourselves.
4.1 Experimental Setup
Evaluation data
We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours (Elazar et al., 2021). We select 15 categories of knowledge triple relations with a larger quantity of data and clearer semantics, totaling 12,671 knowledge triples, as shown in Figure 2. For each category of knowledge triple relations, we create 20 synonymous query templates to construct the QA task evaluation dataset.
Pre-training corpora
  • The Pile: This is an open-source, large-scale text dataset specifically designed for training LLMs. It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality (Gao et al., 2021).
  • ROOTS (En): This dataset is the English subset of ROOTS, a comprehensive multilingual corpus constructed by the BIGScience project. It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency (Laurençon et al., 2022; Scao et al., 2022).
  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs (Soboleva et al., 2023).
  • RefinedWeb: Developed by the Abu Dhabi Technology Innovation Institute (TII), this large-scale English web dataset is based on CommonCrawl and undergoes rigorous content extraction, filtering, and deduplication processes. It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM (Penedo et al., 2023).
  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency (Foundation, ).
Pre-training data retrieval
Pre-trained models
We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs (Vaswani et al., 2017). These models not only make the model weights available but also provide complete pre-training data. They are primarily categorized into the following families:
  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion (Scao et al., 2022).
  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile (Black et al., 2021, 2022; Wang, 2021).
  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters. These models use a uniform data and training process, facilitating research that is both replicable and interpretable (Biderman et al., 2023).
  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata. It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions (Zhang et al., 2024; Soboleva et al., 2023; Li et al., 2023). The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
Evaluation setting
The results of the pre-training data retrieval indicate that the co-occurrence frequency distribution of question entities and answer entities is highly uneven (Figure 2). The majority of QA data co-occur between 10 and 1000 times. In contrast, QA data with higher co-occurrence frequencies are notably sparse. To ensure data balance and effectively evaluate data across different ranges, we group QA data with similar metric values. For co-occurrence metrics and MI metrics, we calculate the average accuracy using intervals of 0.2. The SMI metric is computed based on the MI metric.
@@ 4.1 Experimental Setup
(a) $\text{R}^{2}$ on smaller models
(b) $\text{R}^{2}$ on larger models

### v4

# Beyond Scaling: Measuring and Predicting the Upper Bound of Knowledge Retention in Language Model Pre-Training

@@ Abstract
The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent. Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets. In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task that directly reflects a model’s internalized knowledge without the help of external tools. We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance. In response, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models. We then develop a multi-template QA evaluation framework incorporating paraphrased question variants. Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring additional training. Experimental results show that SMI outperforms co-occurrence-based baselines, achieving $R^{2}>0.75$ on models with over one billion parameters. Theoretical analysis further suggests an upper bound of around 80% QA accuracy under optimal pre-training, reflecting intrinsic memory limitations and motivating the use of retrieval or few-shot methods in later stages.
Abstract
The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent. Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets. In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task that directly reflects a model’s internalized knowledge without the help of external tools. We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance. In response, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models. We then develop a multi-template QA evaluation framework incorporating paraphrased question variants. Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring additional training. Experimental results show that SMI outperforms co-occurrence-based baselines, achieving $R^{2}>0.75$ on models with over one billion parameters. Theoretical analysis further suggests an upper bound of around 80% QA accuracy under optimal pre-training, reflecting intrinsic memory limitations and motivating the use of retrieval or few-shot methods in later stages.
[…]
@@ 1 Introduction
The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" [25]. However, the report provides little detail on the methodology behind this prediction approach.
@@
This technique is critical for the pre-training of large language models (LLMs). On the one hand, the cost of a single pre-training run can reach hundreds of millions of dollars, making re-pre-training a highly challenging endeavor [14, 39, 8]. The prediction technique, which allows for accurate performance forecasting on certain tasks based only on the pre-training signals, helps mitigate the resource waste. On the other hand, it can guide the construction of pre-training data that are more aligned with target tasks, optimizing resource usage [12, 21].
@@
In this paper, we aim to predict performance in closed-book QA, a vital downstream task indicative of a model’s internal knowledge [29]. This setting is particularly well aligned with the nature of pre-training data and best reflects a model’s ability to answer fact-based questions. Meanwhile, we encounter three major challenges (Figure 1):
@@
Challenge1: Limited access to and understanding of pre-training corpora. Currently, few open-source foundation models fully disclose their pre-training datasets. Pre-training from scratch is prohibitively expensive, requiring massive data collection and compute. For example, Meta reports that pre-training the 405B-parameter Llama 3 model used 15.6 trillion tokens and 16,000 H100 GPUs, with estimated costs over 100 million dollars [8].
@@
Challenge2: Limitations of current evaluation methods for pre-trained models. Given the characteristics of closed-book QA tasks, we can evaluate a model’s accuracy on these tasks to assess the factual knowledge it has internalized during pretraining [29, 36]. However, most evaluation methods face challenges, such as being overly sensitive to specific in-context examples and having coarse granularity in test data segmentation [24, 33]. These issues make it difficult to accurately assess whether the pre-trained model has effectively retained specific knowledge. In particular, assessing a model’s knowledge using a single query template is unreliable [40].
@@
Challenge3: Limitations of frequency-based metrics in predicting model performance. The model’s success on the target task relies on its ability to acquire world knowledge during pre-training, a process heavily shaped by the data [1, 37]. Previous research suggests that models tend to acquire knowledge more effectively when it is frequently encountered within the pre-training data [16, 22, 2]. However, these frequency-based methods does not take into account the specificity of knowledge, and thus only demonstrates general trends. For example, despite similar frequencies of occurrence, the model may struggle to pinpoint the location of Apple’s headquarters due to the pervasive nature of the subject. In contrast, identifying NVIDIA’s headquarters might be relatively straightforward. Furthermore, frequency-based methods do not consider the impact of model size.
@@
Second, inspired by the evaluation methods for fine-tuning LLMs [40], we develop a robust multimodal evaluation approach. Recognizing that the same knowledge can correspond to various questioning styles, each style allows for the design of corresponding templates. For each category of question-answering tasks, we design 20 semantically equivalent but syntactically diverse evaluation templates, enabling a more accurate and reliable assessment of the model’s knowledge retention.
Finally, we propose SMI, an information-theoretic metric that linearly associates pre-training data characteristics, model size, and QA accuracy without requiring additional training. Experimental results on 24 pre-trained models indicate that the SMI metric effectively predicts a model’s knowledge retention. Across various model sizes, ranging from 14 million to 176 billion parameters, SMI significantly outperforms the co-occurrence-based baseline. For models with more than 1 billion parameters, the coefficient of determination (R²) between SMI and QA accuracy exceeds 0.75, indicating a strong linear relationship.
Moreover, the comparative analysis across models shows diminishing returns from both scaling model size and optimizing data. As model size increases, its marginal impact on QA performance decreases, suggesting that larger capacity alone cannot guarantee continued gains. Likewise, while higher SMI values are generally associated with better accuracy, pushing SMI further becomes increasingly difficult due to its logarithmic growth. Notably, even in the largest models within each family, QA accuracy consistently plateaus around 80% when SMI is maximized. This suggests a fundamental limitation in how much factual knowledge can be internalized during pre-training, highlighting the need for external augmentation methods such as retrieval or prompting in downstream stages.
In summary, our contributions are threefold:
@@
    Experimental results across diverse models reveal that both scaling model size and optimizing data yield diminishing returns. Even under optimal conditions, QA accuracy plateaus around 80%, reflecting the intrinsic limit of knowledge that can be memorized during pre-training.
@@ 2 Related Work
The characteristics of pre-training data are crucial for knowledge retention and downstream performance in LLMs. Many studies emphasize the role of repetition frequency. For example, Carlini et al. [6] identified a logarithmic relationship between repetition and memory effects, while Chowdhery et al. [7] showed that sequences repeated over 500 times are completed with over 40% accuracy. Ju et al. [15] found that high-frequency data can lead to “factual shortcuts” in multi-hop reasoning. Allen-Zhu and Li [2] suggested that exposing knowledge 1000 times enables a storage capacity of two bits per parameter. Additionally, Razeghi et al. [28] and Yadlowsky et al. [38] linked lower-order co-occurrences to improved numerical reasoning, while McCoy et al. [23] associated task prevalence in pre-training with better performance on tasks like ROT13.
@@
Furthermore, the model size significantly influences the knowledge retention of LLMs. Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern [17].  Kandpal et al. [16] evaluate the BLOOM model families [30] on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters. Additionally, the memory capacity of a model closely correlates with its size [19].
@@
Knowledge triples (subject, relation, object) serve as an effective tool for evaluating the knowledge retention of LLMs. Petroni et al. [27] employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks. Subsequently, He et al. [13], Ju et al. [15] further explore the application of knowledge triples in reasoning tasks. Moreover, Wang et al. [37] and Allen-Zhu and Li [1] use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
@@ 3 Methodology
To address the three challenges outlined in Section 1, we propose a method to predict model knowledge solely from pre-training signals. We begin with prior work and explore the use of the co-occurrence metric. Inspired by information theory, we propose the Mutual Information (MI) metric that simultaneously considers both the frequency and specificity of knowledge, significantly enhancing prediction outcomes. Finally, we propose the SMI metric, a composite representation of pre-training data characteristics and model size.
@@ 3.1 Starting from the Co-occurrence Metric
Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns [16, 22, 2]. Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach [16]. There is a linear relationship between QA accuracy and the logarithm of the number of documents [16]. The logarithm of the document count serves as the co-occurrence metric, which constitutes our baseline.
@@
In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object [37]. For a given LLM, denoted as $F$, we define $F$ as mastering the knowledge triple $t$, thus demonstrating competence in the corresponding QA task, as follows:
@@
Evaluating the knowledge retention of LLMs presents challenges due to the diverse range of queries formed by the subject and relation of knowledge triples. Inspired by the evaluation method for fine-tuned LLMs [40], we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
@@
On the other hand, by employing the multi-template evaluation approach, we measure the model’s accuracy on QA tasks. To examine the relationship between the co-occurrence metric and QA accuracy, we fit a linear regression model and compute the R² along with the Mean Squared Error (MSE). These metrics allow us to evaluate both the strength and the stability of the linear relationship.
@@ 3.2 Considering the Knowledge Specificity to Obtain the MI Metric
The co-occurrence metric does not account for the specificity of knowledge. When entities of knowledge are common phrases that frequently appear in pre-training data, their specificity is low, making the knowledge more difficult to retain. Conversely, if the entities appear infrequently in pre-training data, their specificity is high, and the knowledge is easier to retain. For example, the address of NVIDIA’s headquarters is easier to remember than that of Apple’s headquarters, because the word "Apple" occurs more frequently across diverse contexts. MI in information theory considers both the co-occurrence of two variables and their individual specificities [31]. Therefore, we aim to use MI to improve the co-occurrence metric.
@@
[…]
@@
In the above formula, $P(s)$ and $P(o)$ penalize the number of documents containing the subject and object, respectively, to emphasize the specificity of $s$ and $o$. The term $I(s,o)$ represents the amount of shared information between $s$ and $o$, and it can also indicate the reduction in the amount of information about $o$ given $s$. During inference with LLMs, this amount of information refers to the probability of the model generating the object given the subject. Considering the uneven distribution of $I(s,o)$ (Figure 2), we take its logarithm to obtain the MI metric.
@@ 3.3 Considering the Model Size to Obtain the Size-dependent MI metric
The MI metric does not account for the impact of model size. Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger [17]. In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters [16]. Therefore, we propose to enhance the MI metric by incorporating model size $\Phi$, measured in billions of parameters.
@@
4 Experiments
In the Experimental Setup section, we begin by cleaning and enhancing an open-source Wikipedia knowledge triple dataset to make it suitable for inference with pre-trained models. We then introduce five large-scale open-source pre-training corpora and perform retrieval-based analyses on them. This is followed by a description of five families of pre-trained language models and the corresponding evaluation protocol. In the Experimental Results section, we analyze the performance of 21 open-source pre-trained models along with 3 models that we pre-trained ourselves, comparing them across three evaluation metrics. We also conduct a hyperparameter study to further explore model behavior under different configurations.
4.1 Experimental Setup
Evaluation data
We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours [9]. We select 15 categories of knowledge triple relations with a larger quantity of data and clearer semantics, totaling 12,671 knowledge triples, as shown in Figure 2. For each category of knowledge triple relations, we create 20 synonymous query templates to construct the QA task evaluation dataset.
Pre-training corpora
  • The Pile: This is an open-source, large-scale text dataset specifically designed for training LLMs. It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality [11].
  • ROOTS (En): This dataset is the English subset of ROOTS, a comprehensive multilingual corpus constructed by the BIGScience project. It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency [18, 30].
  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs [32].
  • RefinedWeb: Developed by the Abu Dhabi Technology Innovation Institute (TII), this large-scale English web dataset is based on CommonCrawl and undergoes rigorous content extraction, filtering, and deduplication processes. It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM [26].
  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency [10].
Pre-training data retrieval
Pre-trained models
We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs [34]. These models not only make the model weights available but also provide complete pre-training data. They are primarily categorized into the following families:
  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion [30].
  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile [4, 5, 35].
  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters. These models use a uniform data and training process, facilitating research that is both replicable and interpretable [3].
  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata. It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions [41, 32, 20]. The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
Evaluation setting
The results of the pre-training data retrieval indicate that the co-occurrence frequency distribution of question entities and answer entities is highly uneven (Figure 2). The majority of QA data co-occur between 10 and 1000 times. To ensure data balance and effectively evaluate data across different ranges, we group QA data with similar metric values. For co-occurrence metrics and MI metrics, we calculate the average accuracy using intervals of 0.2. The SMI metric is computed based on the MI metric.